### v1

@@ 4.2.2 Parallel Training
Figure 8: FL-LLaMA: Noise Injection across different number of blocks in $Blocks_{A}$ on MultiRC dataset.
4.2.3 Effects of Noise Scales
@@
The magnitude of noise injection significantly influences the stability and convergence speed of model training. Given that noise is directly injected into the hidden state $h_{A}$, the relative magnitude between the noise and $h_{A}$ naturally affects the model’s performance. To quantify this effect, we first fine-tuned the FL-LLaMA model on the MultiRC [68] with a batch size of 2 and a learning rate of $1\times 10^{-5}$, analyzing the average absolute magnitude of all elements within hidden state $h_{A}$. According to Fig. 7, the average magnitude of the hidden state $h_{A}$ is approximately 0.02 ($num_{Blocks_{A}}$=1). The average magnitude of hidden state $h_{A}$ increases with the number of blocks in $Blocks_{A}$. Based on this observation, we subsequently injected noise at various scales around this value during training to investigate how the noise magnitude affects both the convergence speed and downstream task performance. Training is conducted for 5,000 steps with FP16 precision. We used LoRA with a rank of 8 and a scaling factor $\alpha=16$. The learning rate is 1e-5, and the batch size is 2.
The corresponding experimental results are presented in Fig. 8(a)-8(c) and Tab. 4. Fig. 8(a)-8(c) demonstrate that when the number of blocks in $Blocks_{A}$ and $Blocks_{C}$ are held constant, the convergence speed of FL-LLaMA decreases as the magnitude of injected noise increases. When the average absolute value of the noise is small (e.g., below 0.02), the loss curve remains close to that of the noise-free baseline. A horizontal comparison reveals that, under the same noise level, increasing the number of blcoks in $Blocks_{A}$ consistently accelerates convergence.
@@
Tab. 4 indicates that under fixed noise levels, increasing the number of blocks in $Blocks_{A}$ consistently improves downstream performance in terms of both F1 and EM scores. Notably, when the noise amplitude reaches 0.1, only the single-block configuration of $Blocks_{A}$ shows a significant drop in downstream performance, whereas the three-block version of $Blocks_{A}$ demonstrates greater robustness. This can be explained by the average magnitude shown in Fig. 7: as the depth of $Blocks_{A}$ increases, the ratio between the average absolute value of the hidden state $h_{A}$ and the noise amplitude also increases, thereby improving the signal-to-noise ratio (SNR). This enhanced SNR is precisely what contributes to the better convergence and stronger robustness observed with a larger number of $Blocks_{A}$.
Noise Scale Blocks_A = 1 Blocks_A = 2 Blocks_A = 3
0.0 77.48/40.82 77.82/39.24 77.89/41.34
0.02 77.89/38.09 77.87/39.66 78.61/42.39
0.05 74.57/29.17 77.07/38.41 78.95/41.03
0.1 59.11/8.18 72.45/23.71 75.68/34.10
4.2.4 Model Reverse Attack
To investigate the impact of split point selection and noise scale on the security of FL-LLaMA, we evaluate the attack performance of the adversarial model under varying configurations of $Blocks_{A}$ depth (Blocks_A = 1, 2, 3) and noise (scale = 0, 0.02, 0.5). The attack model is trained via full-parameter fine-tuning with a learning rate of 2e-5. Reconstruction quality is evaluated using standard generation metrics BLEU-4 and ROUGE-2. FedBert reports a Rouge-2 of 7.053 and a Bleu-4 of 28.57. Experiment results of FL-LLaMA is shown in Tab. 5. From the results, we can see that when the client holds only the embedding layer (similar to FedBERT), the attack model achieves relatively high BLEU and ROUGE scores. However, under the FL-LLaMA framework, where the client also retains a portion of the LLMs block in addition to the embedding layer, all attack metrics approach zero. Tab. 6 reports the downstream performance of $client_{0}$ after multi-client training, showing that secret attack allows training to proceed as usual without disrupting the federated process. These findings demonstrate that FL-LLaMA effectively mitigates model inversion attacks, thereby preserving the privacy of client-side data.
Table 5: Security analysis (Rouge-2 F1 / BlEU-4) of FL-LLaMA on Record dataset.
Noise scale Blocks_A=1 Blocks_A=2 Blocks_A=3
0.0 0.051/0.003 0.056/0.003 0.069/0.016
0.02 0.050/0.002 0.052/0.003 0.084/0.044
0.05 0.050/0.003 0.051/0.002 0.067/0.004
Table 6: Performance (EM) under different noise scales and blocks number of $Blocks_{A}$ on Record dataset.
Noise Scale Blocks_A = 1 Blocks_A = 2 Blocks_A = 3
0.0 77.4 79.8 78.9
0.02 79.1 79.9 79.6
0.05 79.6 79.4 79.2
4.2.5 Efficiency Optimization
Figure 9: FL-LLaMA: Inference Speed Comparison (w/ vs. w/o Attention Mask).
Fig. 9 shows a comparison of the time consumption between transferring the attention mask and transferring only the actual input length during distributed training. The bar chart represents the average time spent per token in FL-LLaMA’s forward propagation, while the line chart represents the number of tokens generated per second by FL-LLaMA. It can be observed from the figure that transferring the compressed effective information of the attention mask plays a significant role in reducing the latency of distributed LLMs, with an average time reduction of 87.6%.
Figure 10: FL-LLaMA: Inference Speed Comparison(w/ vs. w/o KV Cache).
Fig. 10 illustrates the effect of the KV cache technique on improving inference speed for different query lengths in distributed inference. The experimental results indicate that without the KV cache, the inference speed of FL-LLaMA significantly decreases as the query length increases, primarily due to the continuously increasing number of K and V vectors computed at each step in the self-attention mechanism. In contrast, with the KV cache technique, the client only needs to send the intermediate vector corresponding to a single token to the server, while reusing the previously generated K and V vectors, thereby effectively avoiding redundant computations. When the query length reaches 4000, the inference speed of FL-LLaMA with KV cache is 8.2 times that without it, and as the query length increases further, this advantage will continue to expand, greatly alleviating the computational burden and operational latency of distributed LLMs.
4.2.6 Adaptive Splitting
The choice of model partitioning has a significant impact on the computational burden at the client side. To evaluate whether FL-LLaMA maintains robust performance under different partitioning strategies, we investigate its downstream task performance when clients hold different number of blocks in $Blocks_{A}$ and $Blocks_{C}$ (Blocks_A = 1, 2, 3 and Blocks_C = 1, 2, 3). This setting reflects realistic constraints where clients possess limited computational resources. All models are trained for 8,000 steps using FP16 precision. LoRA is applied with a rank of 16 and a scaling factor of $\alpha=32$. The training used a batch size of 2 and a learning rate of 2e-5.
Blocks_A Blocks_C = 1 Blocks_C = 2 Blocks_C = 3
1 80.34/44.60 80.63/46.59 80.90/47.11
2 80.41/45.86 80.62/45.96 79.94/45.02
3 81.17/47.22 80.74/46.59 80.20/45.33
Tab. 7 presents F1a/EM scores under various partitioning schemes. The highest F1a/EM scores (81.17/47.22) emerge when Blocks_A=3 and Blocks_C = 1, indicating that allocating more early blocks to the client side can slightly enhance task performance. Overall, the performance differences across partitions remain small, suggesting that FL-LLaMA is relatively insensitive to how the model is split. From a practical standpoint, these findings confirm that clients with constrained resources can flexibly adjust how many blocks they store without substantially degrading accuracy. In scenarios where client-side memory or compute capacities are limited, reducing Blocks_A and Blocks_C remains viable, given that FL-LLaMA maintains robust performance across a range of partitions.
Building upon partitioning experiment, we introduced controlled noise into the hidden states produced by $Blocks_{A}$. Specifically, Gaussian noise with a scale of 0.02 is added during forward propagation before sending the hidden states to the server. Tab. 8 shows that while there is a slight drop compared to the noise-free scenario, the overall performance remains robust, indicating that moderate perturbations in the hidden states do not severely impact convergence or accuracy. Futher, the experimental results show that the higher the number of $Blocks_{A}$, the smaller the performance degradation of FL-LLaMA on downstream tasks. This indicates that the resistance to noise is higher, which verifies that the signal-to-noise ratio has a significant impact on the performance of the model.
Table 8: Different model partitions analysis(F1a/EM) on MultiRC dataset. Noise (scale=0.02) is added to the states $h_{A}$.
Blocks_A Blocks_C = 1 Blocks_C = 2 Blocks_C = 3
1 79.44/38.41 79.58/38.61 79.45/39.66
2 80.10/41.55 80.14/38.61 79.99/40.40
3 79.90/40.29 80.29/43.02 80.44/42.18
5 Conclusion
In this work, we proposed FL-LLaMA, a novel federated split learning framework tailored for large language models. By securely offloading the majority of model computation to a central server while maintaining privacy-sensitive components on local clients, FL-LLaMA addresses the core limitations of traditional federated split learning approaches. Through secure end-to-end propagation with noise injection, parallel training, inference acceleration, and adaptive partitioning tailored to downstream tasks and hardware, FL-LLaMA demonstrates significant advances in security, efficiency, and adaptability. This federated split LLMs framework not only keeps users’ private data confined to their local devices but also significantly shortens training and inference time, making it better suited for handling the scale and complexity of LLMs. In future work, we aim to support more basement models and deploy FL-LLaMA framework in scenarios where data privacy is a critical concern.

### v2

@@ Appendix A Noise Added to Hidden States
Figure 6: FedSEA-LLaMA: Noise Injection across different number of blocks in $Blocks_{A}$ on MultiRC dataset.
@@
The magnitude of noise injection significantly influences the stability and convergence speed of model training. Given that noise is directly injected into the hidden state $h_{A}$, the relative magnitude between the noise and $h_{A}$ naturally affects the model’s performance. To quantify this effect, we first fine-tune the FedSEA-LLaMA model on the MultiRC (Khashabi et al. 2018) with a batch size of 2 and a learning rate of $1\times 10^{-5}$, analyzing the average absolute magnitude of all elements within the hidden state $h_{A}$. According to Figure 7, the average magnitude of the hidden state $h_{A}$ is approximately 0.02 ($num_{Blocks_{A}}$=1). The average magnitude of the hidden state $h_{A}$ increases with the number of blocks in $Blocks_{A}$.
Based on this observation, we subsequently inject noise at various scales around this value during training to investigate how the noise magnitude affects both the convergence speed and downstream task performance. Training is conducted for 5,000 steps with FP16 precision. We use LoRA with a rank of 8 and a scaling factor $\alpha=16$. The learning rate is 1e-5, and the batch size is 2. Corresponding experimental results are presented in Figure 6(a)-6(c) and Table 2.
@@
A horizontal comparison reveals that, under the same noise level, increasing the number of blocks in $Blocks_{A}$ consistently accelerates convergence. Table 2 indicates that when the noise amplitude reaches 0.1, only the single-block configuration of $Blocks_{A}$ shows a significant drop in downstream performance, whereas the three-block version of $Blocks_{A}$ demonstrates greater robustness. This can be explained by the average magnitude shown in Figure 7: as the depth of $Blocks_{A}$ increases, the ratio between the average absolute value of the hidden state $h_{A}$ and the noise amplitude also increases, thereby improving the signal-to-noise ratio (SNR). This enhanced SNR is precisely what contributes to the better convergence and stronger robustness observed with a larger number of $Blocks_{A}$.
Appendix B Prompt
Table 5 illustrates examples of cloze-style questions and corresponding answers derived from each task. (1)ReCoRD tests reading comprehension with common sense reasoning by resolving ambiguous entities in passages. (2)COPA assesses causal reasoning by choosing the more likely cause or effect for a given premise. (3)WSC focuses on pronoun disambiguation requiring deep contextual understanding. (4)RTE evaluates textual entailment by determining if one sentence follows, contradicts, or is unrelated to another. (5)BoolQ involves yes/no questions answered using evidence from context passages. (6)WiC tests word sense disambiguation by checking if a word has the same meaning in two sentences. (7)CB deals with entailment in complex sentences with nuanced reasoning. (8)MultiRC features multi-answer questions based on multi-sentence contexts, requiring integration of dispersed information.
Table 6 illustrates examples of fine-tuning templates of CoQA and XSum.
Table 5: Cloze questions and answers for the 8 SuperGLUE tasks
Dataset Task Fine-tuning template
CoQA conversational QA
<s><s>[INST] <<SYS>><<SYS>>You are a helpful assistant.
You need to answer my question base on the provided story.<</SYS>><</SYS>>
The story is {story}.My question is {question}.[/INST]{answer}</s></s>
XSum summarization generation
<s><s>[INST] <<SYS>><<SYS>>You are a helpful assistant.
You need to summarize an abstract base on the provided passage.<</SYS>><</SYS>>
The passage is {passage}.[/INST]{summarization} </s></s>
Table 6: Fine-tuning templates for CoQA and XSum
Appendix C Parallel Training
num. of clients 2 2 4 8 2 4 8 -
[…]
(a) Client-batch parallel training.
(b) Server-hierachical parallel training.
Figure 8: Multi-client Parallel Training of FedSEA-LLaMA. In (a), the server receives forward-pass states from multiple clients simultaneously, concatenates them into a batch, and performs joint forward propagation. Then, the server splits hidden state $h_{B}$ and sends them back to different clients. In (b), each $client_{i}$ is paired with a corresponding $sub\_server_{i}$. After several training steps, the central server aggregates all sub-servers’ models using FedAvg and distributes the updated global server model back to each $sub\_server_{i}$.
In multi-client settings, the FedSEA-LLaMA framework accommodates three training paradigms: sequential training, client-batch parallelism, and server-hierarchical parallelism, each designed for different system configurations and resource availabilities.
Sequential Training
Client-batch Parallelism
The client-batch parallel approach enhances training throughput by expanding the effective batch size at the server, as shown in Figure  8(a). During each training round, the server collects intermediate hidden states from all participating clients. If each of the $M$ clients processes a local batch of size $b$, the hidden states sent to the server will each have the shape $(b,\textit{seq\_length},\textit{hidden\_size})$. The server concatenates these hidden states along the batch dimension, forming a combined tensor of shape $(b\times M,\textit{seq\_length},\textit{hidden\_size})$. This aggregated batch is then used to perform a joint forward and backward pass on the server model, allowing it to compute gradients more efficiently and update shared parameters.
Server-hierarchical Parallelism
In a server-hierarchical parallel setup, the central server spawns multiple sub-server model instances, each dedicated to a specific client, illustrated in Figure  8(b). These instances run in parallel, enabling simultaneous local training for all clients. Typically, after a fixed number of local updates, the central server synchronizes the training progress by aggregating model parameters across all sub-servers. This is usually done via weighted averaging of both server-side and client-side model parameters. This strategy enables significant improvements in training efficiency and scalability, especially in environments where the server has access to multiple GPUs or multi-threaded execution capabilities.
For communication cost analysis, we randomly select 1,000 samples from the ReCoRD dataset. All training uses a batch size of 1 with FP16 precision for efficiency. LoRA is applied with a rank of 8 and a scaling factor of 16, targeting attention modules during fine-tuning. The model is partitioned with Blocks_A=1 and Blocks_C=1. Each experiment is repeated five times, and we report the mean and standard deviation of the total communication time in Table 7. Results show that sequential training is the slowest, client-batch training is faster, and server hierarchical training is the fastest. Notably, when the server can handle larger batches, increasing the number of clients in client-batch training effectively reduces training time.
Appendix D Parameter Setting
FedSEA-LLaMA and Baselines
Different hyperparameters are explored to obtain the best performance of FedSEA-LLaMA and Centralized LLaMA. Specifically, we tune the LoRA rank (r) and scaling factor (alpha) as pairs, trying combinations including (4, 32), (8, 32), (16, 32), (32, 32), and (8, 16). For the learning rate (lr), values of 1e-5, 2e-5, and 3e-5 are tested. The final parameter settings were selected based on the model’s validation performance. The final parameters are shown in Table 8.
Table 8: Hyperparameters for FedSEA-LLaMA and Baselines
CoQA XSum MultiRC Record
FedSEA-LLaMA Lora r & alpha $(4,32)$ $(16,32)$ $(32,32)$ $(32,32)$
lr 1e-4 3e-5 1e-5 2e-5
Centralized-LLaMA Lora r & alpha $(8,16)$ $(32,32)$ $(32,32)$ $(8,32)$
lr 5e-5 3e-5 1e-5 3e-5
Different Noise Scales and Layers of $Block_{A}$ Experiment
We use a LoRA rank of 8 with a scaling factor of 16, a learning rate of 1e-5, and FP16 precision to investigate the effects of different noise scales and the number of blocks in $Blocks_{A}$ on downstream performance.
Ablation Study on Attention Mask
The context length is set to 500, and the precision is set to FP16.
Ablation Study on KV Cache
The precision is set to FP16. Different context lengths are set to investigate the inference speed of FedSEA-LLaMA and FL-GLM.
Different Model Partition
All experimental results were obtained after 8,000 training steps with LoRA rank set to 8, scaling factor set to 16, a batch size of 1, FP16 precision, and a learning rate of 1e-5.